A can of paint covers 2.5 m2 of a wall. How many cans are needed to paint a square wall with a side length of 6 meters?
1
Expert's answer
2013-04-15T11:30:50-0400
First ofall we need to find the area of the wall. Area of square can be found by the
next formula: S=a*a, where 'a' is aside length of square. So, in our case we have a=6 m, therefore we obtain: S = 6*6=36(m^2)
One can ofpaint covers 2.5m^2 of a wall, sowe can calculate a number of the cans needed as: 36/2.5 = 14.4 (cans).
But wecan’t buy only 0.4 of the one can! So, the number of cans is the closest
greater integer to 14.4 and that is 15.
Answer:15.